# Observable Propagation: A Data-Efficient Approach to Uncover Feature Vectors in Transformers

### v1

@@ C.1 LayerNorm gradients are inversely proportional to input norms
In §2.4, it was stated that LayerNorm gradients are not constant, but instead, depend on the norm of the input to the LayerNorm. To elaborate, the gradient of $n^{T}(\sqrt{d}W\operatorname{LayerNorm}(x)+b)$ can be shown to be $\frac{\sqrt{d}W}{\left\|Px\right\|}P\left(I-\frac{(Px)(Px)^{T}}{\left\|Px\right\|^{2}}\right)n$ (see Appendix H). $P$ and $\left(I-\frac{(Px)(Px)^{T}}{\left\|Px\right\|^{2}}\right)$ are both orthogonal projections that leave $\left\|n\right\|$ relatively untouched, so the term that is most responsible for affecting the norm of the feature vector is the $\frac{\sqrt{d}W}{\left\|Px\right\|}$ factor. Now, by Lemma 1 in Appendix H, we have that $\frac{\sqrt{d}W}{\left\|Px\right\|}\approx\frac{\sqrt{d}W}{\left\|x\right\|}$. Thus, if $\widetilde{\left\|x\right\|}$ a good estimate of $\left\|x\right\|$ for a given set of input prompts at a given layer, then a good approximation of the gradient of a LayerNorm sublayer is given by $\left(\sqrt{d}W/\widetilde{\left\|x\right\|}\right)n$. This approximation can be used to speed up the computation of gradients for LayerNorms.
C.2 Feature vector norms with LayerNorms
@@
In §4.1, we explained that looking at the norms of feature vectors can provide a fast and reasonable guess of which model components will be the most important for a given task. However, there is a caveat that must be taken into account regarding LayerNorms. As shown in Appendix C.1, the gradient of a LayerNorm sublayer is approximately inversely proportional to the norm of the input to the LayerNorm sublayer.
@@
Now, assume that we have a computational path beginning at a LayerNorm, where $\widetilde{\left\|x\right\|}$ is an estimate of the norm of the inputs to that LayerNorm. Let $y$ be the feature vector for this computational path. Then we have $y\approx\sqrt{d}W/\widetilde{\left\|x\right\|}y^{\prime}$, where $y^{\prime}$ is the feature vector for the “tail” of the computational path, that comes after the initial LayerNorm.
@@
Appendix D Details on linear approximations for MLPs
Thus, we have two points $x_{-}$ and $x_{+}$ such that $g(x_{-})<0$ and $g(x_{+})>0$. Since MLPs are continuous, there therefore must be some point $x^{*}$ at which $g(x^{*})=0$: a point that lies on the decision boundary of the MLP. It stands to reason that the gradient at this decision boundary is more likely to capture the larger-scale behavior of the MLP and is less likely to be saturated, when compared to the gradient at more “extreme” points like $x_{-}$ and $x_{+}$.
Appendix E Empirical LayerNorm gradient investigations
In this section, we put forth various empirical results relevant for the discussion of LayerNorm gradients in §2.4.
E.1 LayerNorm input norms per layer
We calculated the average norms of inputs to each LayerNorm sublayer in the model, over the activations obtained from ten of the prompts from the artificial dataset described in Appendix B. The results can be found in Figure 3. The wide variation in the input norms across different layers implies that input norms must be taken into account in any approximation of LayerNorm gradients.
Figure 3: Mean norms of activations before each LayerNorm
E.2 LayerNorm weight values are very similar
In §2.4, we state that the entries in the LayerNorm scaling matrices tend to be very close together, and use this as justification for treating weight matrices as scalars. Specifically, we found that the average variance of scaling matrix entries across all LayerNorms in GPT-Neo-1.3B is $0.007827$. To determine the extent to which this variance is large, we calculated the ratio of the variance of each LayerNorm’s weight matrix’s entries to the mean absolute value of each layer’s embeddings’ entries. The results can be found in Figure 4. Note that the highest value found was $0.0731$ at Layer 0 – meaning that the average entry in Layer 0 embeddings was over 13.67 times larger than the variance between entries in that layer’s ln_1 LayerNorm weight. This supports our assertion that LayerNorm scaling matrices can be largely treated as constants.
One possible guess as to why this behavior might be occurring is this: much of the computation taking place in the model does not occur with respect to basis directions in activation space. However, the diagonal LayerNorm weight matrices can only act on these very basis directions. Therefore, the weight matrices end up “settling” on the same nearly-constant value in all entries.
Appendix F Further debiasing experiments
We ran further experiments on the artificial dataset described in Appendix B, in order to determine the extent to which the feature vectors yielded by observable propagation could be used for debiasing the model’s outputs. The idea is similar to that presented by Li et al. (2023): by adding a feature vector to the activations at a given layer for the name token, we can hopefully shift the model’s output to be less biased.
In particular, let $y$ be the unit 6::6 feature vector for $n_{\text{bias}}$, let $x_{\text{female}}$ be the activation vector for the name token at that layer for the female prompt, and let $x_{\text{male}}$ be the activation vector for the name token at that layer for the male prompt. Then we added the vector $y^{\prime}=\left((x_{\text{female}}-x_{\text{male}})\cdot y\right)y$ to $x_{\text{male}}$. If the model were a linear model whose output was solely determined by the dot product of the input at this layer with the feature vector $y$, then the output of the model in the case where $y^{\prime}$ is added to the male embeddings would be the same as the output of the model on the female prompt. Therefore, the difference between this “patched” output and the model’s output on the female prompt can be viewed as an indicator of the extent to which the feature vector is affected by nonlinearity in the model. We also ran this same experiment, but adding $2y^{\prime}$ instead of $y^{\prime}$ to the male embeddings, in order to get a stronger debiasing effect.
Mean $n_{\text{bias}}$ score
$-2.947$ $5.011$ $-0.5489$ $4.879$
Mean absolute difference with female scores
$7.9903$ 00 $5.6245$ $1.3180$
Mean difference from female scores
$7.9583$ 00 $5.5599$ $0.1316$
Table 4: The results of the debiasing experiments for the $n_{\text{bias}}$ observable. “Mean absolute difference with female scores” refers to the mean absolute difference between the $n_{\text{bias}}$ score for each male prompt (or male prompt with patched activations) and the score for the corresponding female prompt. “Mean difference from female scores” refers to the mean difference, without taking the absolute value, between the $n_{\text{bias}}$ for the female prompt, and the score for the corresponding (patched) male prompt.
We then wanted to investigate the extent to which adding this “debiasing vector” would harm the model’s performance on the pronoun prediction task. As such, we repeated these experiments on the dataset of prompts for $n_{\text{subj}}$, but adding $2y^{\prime}$ to the male activations. The results can be found in Table 5. The results show that adding the “debiasing vector” to the male name embeddings also causes the model’s ability to correctly predict gendered pronouns to drop dramatically. This suggests that in cases such as this one, where the model uses the same features for undesirable outputs as it does for desirable outputs, inference-time interventions such as that presented by Li et al. (2023) may cause an inevitable decrease in model quality.
Male prompt Female prompt Male patched $+2y^{\prime}$
Mean $n_{\text{subj}}$ score
$-5.1393$ $5.0404$ $4.8794$
Mean absolute difference with female scores
$10.180$ 00 $2.148$
Mean difference from female scores
$10.180$ 00 $0.161$
Appendix G Experimental details for Section 4.3
G.1 Datasets
The dataset used in the subject pronoun prediction task is the same artificial dataset described in Appendix B.
The dataset used in the American political party prediction task is an artificial dataset consisting of prompts of the form "[NAME] is a", where [NAME] is replaced by the name of a politician drawn from a list of 40 Democratic Party politicians and 40 Republican Party politicians. These politicians were chosen according to the list of “the most famous Democrats” and “the most famous Republicans” for Q3 2023 compiled by YouGov, available at https://today.yougov.com/ratings/politics/fame/Democrats/all and https://today.yougov.com/ratings/politics/fame/Democrats/all. The intuition behind this choice of dataset is that the model would be more likely to identify the political affiliation of well-known politicians, because better-known politicians would be more likely to occur in its training data. This is the primary reason that a smaller dataset is being used.
G.2 Task definition
G.3 Feature vectors
G.4 Task evaluation
Appendix H Proof of Theorem 1
@@
Using the multivariate chain rule along with the rule that the derivative of $\frac{x}{\|x\|}$ is given by $\frac{I}{\|x\|}-\frac{xx^{T}}{\|x\|^{3}}$ (see §2.6.1 of Petersen & Pedersen (2012)), we thus have that
@@
Appendix I Empirical results regarding Theorem 1
Note that Theorem 1 assumes that feature vectors $n$ are normally-distributed, which may not necessarily occur in practice (although given that observables and observable-derived feature vectors are only introduced in this work, it is hard to say whether this is false more generally; more research is necessary, including research on the sorts of observables that practitioners wish to analyze in practice). However, the intention of Theorem 1 is to provide motivation that underpins what we found empirically: namely, that feature vectors computer by taking LayerNorm into account have extremely high cosine similarities with feature vectors computed without taking LayerNorm into account.
In particular, for the feature vectors considered in Section 4.3, these cosine similarities and angles are given in Table 6. For reference, note that the upper bound on the angle between these feature vectors according to Theorem 1 is approximately 0.0442 radians. The feature vectors for subject pronoun prediction have a higher angle between them of 0.0664 radians, but this can be attributed to the fact that the circuit for these feature vectors goes through multiple LayerNorms. Additionally, the angle for the political party prediction feature vector is also slightly higher than the bound predicted by the theorem; but it is worth noting that the theorem predicts a bound on the expected angle, rather than a bound on the maximum angle; this also might be due to the scaling matrix $W$ in the LayerNorm (see Appendix E).
@@ Appendix I Empirical results regarding Theorem 1
Appendix J Proof of Theorem 2
@@
Appendix K Top activating tokens on 1M tokens from The Pile for $n_{\text{bias}}$ and $n_{\text{subj}}$ 6::6 feature vectors
In §4.2, in order to confirm that the feature vectors that we found for attention head 6::6 corresponded to notions of gender, we looked at the tokens from a dataset of 1M tokens from The Pile (see Appendix B) that maximally and minimally activated these feature vectors.
K.1 $n_{\text{bias}}$ feature vector
@@
K.2 $n_{\text{subj}}$ feature vector

### v2

@@ E.3 LayerNorm Gradients Are Inversely Proportional to Input Norms
In §2.4, it was stated that LayerNorm gradients are not constant, but instead, depend on the norm of the input to the LayerNorm. To elaborate, the gradient of $n^{T}(\sqrt{d}W\operatorname{LayerNorm}(x)+b)$ can be shown to be $\frac{\sqrt{d}W}{\left\|Px\right\|}P\left(I-\frac{(Px)(Px)^{T}}{\left\|Px\right\|^{2}}\right)n$ (see Appendix E.5). $P$ and $\left(I-\frac{(Px)(Px)^{T}}{\left\|Px\right\|^{2}}\right)$ are both orthogonal projections that leave $\left\|n\right\|$ relatively untouched, so the term that is most responsible for affecting the norm of the feature vector is the $\frac{\sqrt{d}W}{\left\|Px\right\|}$ factor. Now, by Lemma 1 in Appendix E.5, we have that $\frac{\sqrt{d}W}{\left\|Px\right\|}\approx\frac{\sqrt{d}W}{\left\|x\right\|}$. Thus, if $\widetilde{\left\|x\right\|}$ a good estimate of $\left\|x\right\|$ for a given set of input prompts at a given layer, then a good approximation of the gradient of a LayerNorm sublayer is given by $\left(\sqrt{d}W/\widetilde{\left\|x\right\|}\right)n$. This approximation can be used to speed up the computation of gradients for LayerNorms.
E.4 Feature Vector Norms with LayerNorms
@@
In §4.1, we explained that looking at the norms of feature vectors can provide a fast and reasonable guess of which model components will be the most important for a given task. However, there is a caveat that must be taken into account regarding LayerNorms. As shown in Appendix E.3, the gradient of a LayerNorm sublayer is approximately inversely proportional to the norm of the input to the LayerNorm sublayer.
@@
Now, assume that we have a computational path beginning at a LayerNorm, where $\widetilde{\left\|x\right\|}$ is an estimate of the norm of the inputs to that LayerNorm. Let $y$ be the feature vector for this computational path. Then we have $y\approx\sqrt{d}W/\widetilde{\left\|x\right\|}y^{\prime}$, where $y^{\prime}$ is the feature vector for the ?tail? of the computational path, that comes after the initial LayerNorm.
@@
E.5 Proof of Theorem 1
@@
Using the multivariate chain rule along with the rule that the derivative of $\frac{x}{\|x\|}$ is given by $\frac{I}{\|x\|}-\frac{xx^{T}}{\|x\|^{3}}$ (see §2.6.1 of (Petersen & Pedersen, 2012)), we thus have that
@@
E.6 Empirical Results Regarding LayerNorm Gradients
In Section 2.4, we mention that we empirically found that feature vectors computed by differentiating through LayerNorms had high cosine similarities with feature vectors computed while ignoring LayerNorms.
In particular, for the feature vectors considered in Section 4.3, these cosine similarities and angles are given in Table 4.
@@ E.6 Empirical Results Regarding LayerNorm Gradients
Theorem 1 can be used to estimate the expected angle in radians between a feature vector computed with LayerNorm and a feature vector computed without LayerNorm. In this case, given that the model has dimensionality 512, this expected angle is approximately 0.0442 radians.
In general, this value is a decent estimation to the empirical values that we found – especially when considering that Theorem 1 makes the assumption that feature vectors are normally distributed, and when considering that this theorem does not take into account the scaling matrix after the LayerNorm described in Appendix E.
Additionally, while the feature vectors for subject pronoun prediction have a higher angle between them of 0.0664 radians, this can be possibly be attributed to the fact that the circuit for these feature vectors goes through multiple LayerNorms.
Appendix F Applicability to Modified Transformer Architectures
Many modern transformers include architectural modifications from older transformer models such as GPT2 (Radford et al., 2019). For instance, the open-source Llama 3 family of models (AI@Meta, 2024) modifies attention sublayers by using grouped-query attention (GQA) (Ainslie et al., 2023), and uses RMSNorm (Zhang & Sennrich, 2019) instead of LayerNorm. It is thus natural to consider the extent to which ObProp generalizes to models incorporating these architectural changes. Happily, ObProp functions the exact same way in the presence of GQA and RMSNorm. Because GQA only affects the QK circuit of attention, and ObProp only addresses the OV circuit of attention, GQA does not change the operation of our method.
As for RMSNorm, from a practical perspective, Algorithm 1 still works the same (because RMSNorm can just be treated as another nonlinearity). And from a theoretical perspective, RMSNorm is actually easier to handle than vanilla LayerNorm: for Theorem 1, the bound becomes tighter ($\arccos(\sqrt{1-1/d})$), the normality assumption can be changed to apply to $x$ instead of $n$, and the proof follows almost immediately from Lemma 1. (This is because RMSNorm does not include the $P$ projection found in the proof of Theorem 1.)
Appendix G Further Debiasing Experiments
We ran further experiments on the artificial dataset described in Appendix A, in order to determine the extent to which the feature vectors yielded by observable propagation could be used for debiasing the model’s outputs. The idea is similar to that presented by Li et al. (2023): by adding a feature vector to the activations at a given layer for the name token, we can hopefully shift the model’s output to be less biased.
In particular, let $y$ be the unit 6::6 feature vector for $n_{\text{bias}}$, let $x_{\text{female}}$ be the activation vector for the name token at that layer for the female prompt, and let $x_{\text{male}}$ be the activation vector for the name token at that layer for the male prompt. Then we added the vector $y^{\prime}=\left((x_{\text{female}}-x_{\text{male}})\cdot y\right)y$ to $x_{\text{male}}$. If the model were a linear model whose output was solely determined by the dot product of the input at this layer with the feature vector $y$, then the output of the model in the case where $y^{\prime}$ is added to the male embeddings would be the same as the output of the model on the female prompt. Therefore, the difference between this ?patched? output and the model’s output on the female prompt can be viewed as an indicator of the extent to which the feature vector is affected by nonlinearity in the model. We also ran this same experiment, but adding $2y^{\prime}$ instead of $y^{\prime}$ to the male embeddings, in order to get a stronger debiasing effect.
Mean $n_{\text{bias}}$ score $-2.947$ $5.011$ $-0.5489$ $4.879$
Mean absolute difference with female scores $7.9903$ 00 $5.6245$ $1.3180$
Mean difference from female scores $7.9583$ 00 $5.5599$ $0.1316$
Table 5: The results of the debiasing experiments for the $n_{\text{bias}}$ observable. ?Mean absolute difference with female scores? refers to the mean absolute difference between the $n_{\text{bias}}$ score for each male prompt (or male prompt with patched activations) and the score for the corresponding female prompt. ?Mean difference from female scores? refers to the mean difference, without taking the absolute value, between the $n_{\text{bias}}$ for the female prompt, and the score for the corresponding (patched) male prompt.
We then wanted to investigate the extent to which adding this ?debiasing vector? would harm the model’s performance on the pronoun prediction task. As such, we repeated these experiments on the dataset of prompts for $n_{\text{subj}}$, but adding $2y^{\prime}$ to the male activations. The results can be found in Table 6. The results show that adding the ?debiasing vector? to the male name embeddings also causes the model’s ability to correctly predict gendered pronouns to drop dramatically. This suggests that in cases such as this one, where the model uses the same features for undesirable outputs as it does for desirable outputs, inference-time interventions such as that presented by Li et al. (2023) may cause an inevitable decrease in model quality.
Male prompt Female prompt Male patched $+2y^{\prime}$
Mean $n_{\text{subj}}$ score $-5.1393$ $5.0404$ $4.8794$
Mean absolute difference with female scores $10.180$ 00 $2.148$
Mean difference from female scores $10.180$ 00 $0.161$
Appendix H Proof of Theorem 2
@@
Appendix I Top Activating Tokens on 1M Tokens from The Pile for $n_{\text{bias}}$ and $n_{\text{subj}}$ 6::6 Feature Vectors
In §4.2, in order to confirm that the feature vectors that we found for attention head 6::6 corresponded to notions of gender, we looked at the tokens from a dataset of 1M tokens from The Pile (see Appendix A) that maximally and minimally activated these feature vectors.
I.1 $n_{\text{bias}}$ Feature Vector
@@
I.2 $n_{\text{subj}}$ Feature Vector